OpenStudy (anonymous):

@slaw could use your help again, its more theoretical this time. I have the parametrization but I need help with the domain again. Find the surface area of sphere x^2+y^2+z^2=b^2 above x^2 +y^2 = a^2 where 0

OpenStudy (anonymous):

I got the parametrization \[r(\phi,\theta) = <bsin \phi \cos \theta, bsin \phi \cos \theta, bcos \phi>\]

OpenStudy (anonymous):

well he signed off, can anyone else help me out here?

OpenStudy (anonymous):

Hold on. Be patient

OpenStudy (anonymous):

I am, just wanted to let you guys know he signed off, thats all. My apologies.

OpenStudy (anonymous):

\[ \int _0^{2 \pi }\int _0^ub^2 \sin (\phi )d\phi d\phi =-2 \pi b^2 (\cos (u)-1) \] Where u is the angle that you obtain by joining one of the intersection point of the sphere and the cylinder to the center and the z-axis

OpenStudy (anonymous):

\[ \cos (u)=\frac{\sqrt{b^2-a^2}}{a} \]

OpenStudy (anonymous):

thats with \[d \phi d \theta\] right?

OpenStudy (anonymous):

yes, sorry

OpenStudy (anonymous):

Wait a minute.

OpenStudy (anonymous):

no worries, I appreciate the help. I'll wait

OpenStudy (anonymous):

That is fine the answer is still the same it should be \[ \int _0^{2 \pi }\int _0^ub^2 \sin (\phi )d\phi d\theta =-2 \pi b^2 (\cos (u)-1) \]

OpenStudy (anonymous):

cos(u) is computed above.

OpenStudy (anonymous):

may I ask why it is from 0 to u and not from u to pi/2?, I thought it would be since it was above the cylinder.

OpenStudy (anonymous):

Final Answer must look like \[ \int _0^{2 \pi }\int _0^ub^2 \sin (\phi )d\phi d\theta =2 \pi b^2 \left(1-\frac{\sqrt{b^2-a^2}}{a}\right) \]

OpenStudy (anonymous):

If it is from 0 to pi, you get the whole sphere. To get the area above the cylinder, you should let \(\phi \) varies from 0 to the angle where the cylinder touches the sphere.

OpenStudy (anonymous):

That is the angle u

OpenStudy (anonymous):

Do you understand?

OpenStudy (anonymous):

I'm just trying to figure out the way you find cos(u), what did you look at? the point?

OpenStudy (anonymous):

I am drawing a picture for you. Hold on.

OpenStudy (anonymous):

I think I had my clyinder around the wrong axis. Thanks.

OpenStudy (anonymous):

I understand now

OpenStudy (anonymous):

YW